# LLM-Meta-SR: Learning to Evolve Selection Operators for Symbolic Regression

### v1

@@ B.1.1 Static Analysis
Code 4: Omni Selection
@@ B.1.2 Benchmark Analysis
Test $R^{2}$ Scores: The test $R^{2}$ scores demonstrate that the Omni selection operator achieves an advantage at the beginning of the search, indicating that it is an effective selection operator for anytime performance [68].
@@
B.2 Code Evolved by LLM-Meta-SR without Domain Knowledge
In this section, we analyze the code evolved by LLM-Meta-SR without domain knowledge to understand how an LLM can generate algorithms based solely on its internal knowledge. The evolved code is shown in LABEL:lst:omni-selection-free. We refer to the resulting algorithm as Omni-Zero, as it evolves the model from scratch, similar to AutoML-Zero [real2020automl].
B.2.1 Static Analysis
@@
This cosine-distance-based novelty score is better suited for symbolic regression than Euclidean-distance-based novelty scores, as it is more robust to outliers and cannot be easily manipulated by increasing the distance with some outlier, as discussed in Section B.1.2.
@@
Code 5: Omni Selection Zero
B.2.2 Benchmark Analysis
The Omni-Zero operator is evaluated on the SRBench datasets [la2021contemporary], and the results are shown in Figure 20. As shown by the results, the proposed method achieves medium performance among the benchmarked algorithms. It performs significantly better than tournament selection, as indicated by the p-values in Figure 20. However, compared to selection operators that consider specificity, such as the lexicase selection operator, the evolved selection operator is inferior.
@@ B.2.2 Benchmark Analysis
B.3 A Repaired Version of Omni Selection for Small Datasets
B.3.1 Problem Analysis
@@ B.3.3 Benchmark Analysis
Figure 23: Test $R^{2}$ scores of different selection operators on symbolic regression benchmarks, including the repaired omni selection.
Figure 24: Tree sizes of different selection operators on symbolic regression benchmarks, including the repaired omni selection.
@@
Code 6: The Repaired Version of Omni Selection (Omni-R)
Appendix C Test Cases for Synthetic Evaluation
@@
Appendix D Evolutionary Symbolic Regression
D.1 Solution Representation
The symbolic regression model is represented as a symbolic expression tree with linear scaling [nadizar2024geometric]. For a given expression $\Phi$, the prediction is computed as $\hat{Y}=\alpha\cdot\Phi(X)+\beta$, where the coefficients $\alpha\in\mathbb{R}$ and $\beta\in\mathbb{R}$ are fitted using ridge regression by minimizing the regularized squared error between $\hat{Y}$ and the targets $Y$.
D.2 Algorithm Workflow
The symbolic regression algorithm is implemented using the genetic programming framework [banzhaf1998genetic]. Let $\mathcal{F}^{(t)}=\{\Phi_{1},\Phi_{2},\dots,\Phi_{\lambda}\}$ denote the population of $\lambda$ symbolic expressions at generation $t$. The evolutionary process proceeds as follows:
  • Population Initialization: The initial population $\mathcal{F}^{(0)}$ is generated using the ramped half-and-half method [banzhaf1998genetic].
[…]
D.3 Parameter Settings
The parameters for symbolic regression follow conventional settings. The population size and number of generations are set to match those used in D-Split [imai2024minimum]. To prevent division-by-zero errors, we use the analytical quotient ($\mathrm{AQ}(x,y)=\frac{x}{\sqrt{1+y^{2}}}$[ni2012use] in place of the standard division operator.
Table 3: Parameter settings
Parameter Value
Population size 100
Maximum generations 100
Maximum tree depth 10
Initialization method Ramped half-and-half (depth 0-6)
Function set +,,×$+,\,-,\,\times,\,\mathrm{AQ},\,\sqrt{\cdot},\,\log(1+|\cdot|),\,|\cdot|,\,(\cdot)^{2},$ $\sin_{\pi}(\cdot),\,\cos_{\pi}(\cdot),\,\mathrm{Max},\,\mathrm{Min},\,\mathrm{Neg}$
Crossover rate 0.9
Mutation rate 0.1
D.4 Evaluation Protocol in Meta-Evolution
In the evaluation of selection operators generated by the LLM, all other algorithmic components—including solution initialization, evaluation, generation, and elitism—follow standard practices commonly used in evolutionary symbolic regression. For each dataset $\mathcal{D}_{j}$, the data is split into training and validation subsets with an 80:20 ratio [la2021contemporary]. The regression model is optimized to maximize the training $R^{2}$ score, while the validation $R^{2}$ score is used as the final score of the selection operator. This encourages the selection of operators with implicit regularization capabilities.
D.5 Evaluation Protocol for Discovered Operators
For the experiment evaluating the evolved selection operator on SRBench, the dataset is randomly split into a training set and a test set with an 80:20 ratio [la2021contemporary]. Subsampling is applied to limit the maximum number of training instances to 10,000, consistent with the SRBench protocol [la2021contemporary]. Symbolic regression is run on the training set and evaluated on the test set. All features are standardized, as this has been shown to be beneficial for symbolic regression [owen2022standardization]. Standardization parameters are learned from the training data and applied to both the training and test sets. All symbolic regression experiments are conducted on an AMD Milan 7713 CPU.
Appendix E Baseline Selection Operators
The hyperparameters for the baseline selection operators follow the default settings specified in their respective original papers. For RDS-Tour, the sampling ratio is set to 10% and the tournament size to 7 [geiger2025tournament]. For PLex selection, the temperature parameter is set to 1.0 [ding2023probabilistic]. For DALex, a particularity pressure of 3 is used [ni2024dalex]. For CPS, the first parent is selected via tournament, with the tournament size set to 7 [xu2022genetic].
For Boltzmann sampling with temperature scheduling [shojaee2025llmsr, romera2024mathematical], a temperature decay rule is applied to balance exploration and exploitation over time, similar to the approaches used in LLM-SR [shojaee2025llmsr] and FunSearch [romera2024mathematical]. The sampling probability for individual $i$ is defined as:
@@ Appendix E Baseline Selection Operators
where $s_{i}$ denotes the score of individual $i$, $\tau_{0}=0.1$ is the initial temperature [shojaee2025llmsr], $t$ is the current generation, and $T_{\max}$ is the total number of generations. This scheduling rule linearly decays the temperature over generations and gradually shifts the selection pressure from exploration (high temperature) to exploitation (low temperature) as evolution progresses in evolutionary symbolic regression.
Appendix F Further Investigation on Crossover and Mutation Ratio
@@
Appendix G Analysis of Token Count of Generated Code
Appendix H Analysis on State-of-the-Art Symbolic Regression Algorithms
H.1 Experimental Settings
In this section, we apply the evolved operator to a state-of-the-art symbolic regression algorithm, namely retrieval-augmentation-generation-based symbolic regression (RAG-SR) [zhang2025ragsr], to evaluate the effectiveness of the proposed operator in the context of modern symbolic regression. The only modification made is the replacement of the automatic epsilon lexicase selection [la2019probabilistic] with the repaired Omni selection operator described in LABEL:lst:omni-selection-repaired. All other components remain unchanged. The resulting algorithm is referred to as RAG-SR-Omni. For the experimental datasets, following the setup in Section 4, we use 116 out of 120 regression problems, excluding the four used during meta-evolution.
H.2 Experimental Results
The results are presented in Figure 28, with the top-5 performing algorithms shown separately in Figure 29 for better clarity. These figures demonstrate that RAG-SR-Omni outperforms RAG-SR in terms of median $R^{2}$ scores, model sizes, and training time. Furthermore, as shown by the Pareto front of the ranks of test $R^{2}$ scores and model sizes across different algorithms in Figure 30, RAG-SR-Omni establishes a new Pareto front on SRBench. Collectively, these results support the effectiveness of the LLM-evolved selection operator.
Figure 28: Median $R^{2}$ scores, model sizes, and training time of 26 algorithms on the symbolic regression benchmark [la2021contemporary].
Figure 30: Pareto front of the rank of test $R^{2}$ scores and model sizes for different algorithms.
Appendix I Licenses for used assets
The licenses for the used assets are shown in Table 5.
Table 5: Used assets and their licenses.
Type Asset License Usage
Code DEAP [fortin2012deap] LGPL-3.0 License Evaluation
Code CodeBLEU [ren2020codebleu] MIT License Evaluation
Dataset SRBench [la2021contemporary] GPL-3.0 License Testing
@@ Appendix I Licenses for used assets
In recent years, lexicase selection has shown strong empirical performance and is widely adopted in modern symbolic regression systems [cava2018learning, zhang2025ragsr]. Inspired by lexicase selection, several variants have been developed, including probabilistic lexicase selection [ding2023probabilistic], lexicase-like selection via diverse aggregation (DALex) [ni2024dalex], $\epsilon$-lexicase selection with dynamic split (D-Split) [imai2024minimum], down-sampled lexicase selection [boldi2024informed], and random down-sampled tournament selection [geigerperformance]. These methods share the core idea of emphasizing performance on subsets of training instances to encourage specialization, and they demonstrate advantages in various scenarios [geigerperformance]. However, these operators are manually designed and do not fully satisfy all the desired properties outlined in Section 3.5, highlighting the need for further exploration of selection operators.
@@
Before the LLM era, genetic programming had been widely used for automated evolutionary algorithm design, including the design of fitness functions [fong2024metasr], parameter adaptation techniques [stanovov2022automatic], and update strategies [lones2021evolving]. More recently, reinforcement learning has also been applied to designing update rules in differential evolution [chen2024symbol]. However, these approaches are typically limited to generating only small code segments, as exploring large program spaces remains challenging. The emergence of LLMs has made automated algorithm design more feasible. For example, LLMs have been used to design a differential evolution algorithm competitive with state-of-the-art continuous optimization methods [van2024llamea]. Similarly, LLMs have been employed to improve the self-organizing migrating algorithm [pluhacek2024using]. This growing interest in LLMs motivates the development of efficient LLM-based approaches for automated algorithm design, where semantic awareness and code bloat are two key issues that remain underexplored and warrant further investigation.
@@
LLMs have been employed as optimization tools across various domains [meyerson2024language, song2024position], including neural architecture search [chen2023evoprompting], graph neural network design [mo2025autosgnn], reward function design [maeureka], the traveling salesman problem [liu2024large], prompt optimization [yanglarge], automated machine learning [chan2024mle], solver code generation [JiangShu2025llmopt], symbolic regression [shojaee2025llmsr], and feature engineering [gong2025evolutionary]. Evolutionary search has gained considerable attention in the LLM era due to its effectiveness over random sampling [zhang2024understanding]. Various efforts have aimed to enhance LLM-based evolutionary search, including group relative policy optimization [huang2025calm], direct preference optimization [vsurinaalgorithm], core abstraction prompting [wu2025efficient], quality-uncertainty optimization [chen2024qube], and bi-dimensional co-evolution [zhao2025understanding].

### v2

@@ C.1 Code Evolved by LLM-Meta-SR
Code 1: Omni Selection
@@
Test $R^{2}$ Scores: The test $R^{2}$ scores demonstrate that the Omni selection operator achieves an advantage at the beginning of the search, indicating that it is an effective selection operator for anytime performance (Ye et al. 2022).
@@
C.2 Code Evolved by LLM-Meta-SR without Domain Knowledge
In this section, we analyze the code evolved by LLM-Meta-SR without domain knowledge to understand how an LLM can generate algorithms based solely on its internal knowledge. The evolved code is shown in LABEL:lst:omni-selection-free. We refer to the resulting algorithm as Omni-Zero, as it evolves the model from scratch, similar to AutoML-Zero (Real et al. 2020).
Static Analysis
@@ C.2 Code Evolved by LLM-Meta-SR without Domain Knowledge
This cosine-distance-based novelty score is better suited for symbolic regression than Euclidean-distance-based novelty scores, as it is more robust to outliers and cannot be easily manipulated by increasing the distance with some outlier, as discussed in Section C.1.
@@
Code 2: Omni Selection Zero
Benchmark Analysis
The Omni-Zero operator is evaluated on the SRBench datasets (La Cava et al. 2021), and the results are shown in Figure 17. As shown by the results, the proposed method achieves medium performance among the benchmarked algorithms. It performs significantly better than tournament selection, as indicated by the p-values in Figure 17. However, compared to selection operators that consider specificity, such as the lexicase selection operator, the evolved selection operator is inferior.
@@
C.3 A Repaired Version of Omni Selection for Small Datasets
Problem Analysis
@@ C.3 A Repaired Version of Omni Selection for Small Datasets
Figure 20: Test $R^{2}$ scores of different selection operators on small-scale symbolic regression benchmarks, including the repaired omni selection.
Figure 21: Tree sizes of different selection operators on small-scale symbolic regression benchmarks, including the repaired omni selection.
@@
Code 3: The Repaired Version of Omni Selection (Omni-R)
Appendix D Further Analysis of the Semantic-Aware Evolution
In this paper, we propose a semantic-aware evolutionary approach that explicitly selects complementary pairs of individuals for crossover. To illustrate how two complementary codes work together to generate improved solutions, we provide a concrete example in this section. In the existing literature, the most common approach to combining the capabilities of two selection operators is to embed them within a dynamic algorithm selection framework (Xue et al. 2022), allowing each operator to be applied under different conditions. While this strategy permits both operators to contribute, it does so sequentially and without integrating their underlying semantics. Moreover, repeatedly combining two code fragments over multiple generations can lead to exponential code growth (Martins et al. 2018), impairing interpretability and maintainability. In contrast, our approach enables a more meaningful integration by combining semantically complementary code fragments through crossover. As shown in LABEL:lst:offspring and its parent codes in LABEL:lst:parent_a and LABEL:lst:parent_b, the resulting offspring code fragments are not simply concatenations but structured combinations of meaningful components from both parents, as detailed in Table 4. Compared to the mutation operator, which performs a random search around a single solution, crossover provides a directional search by explicitly fusing different capabilities. Thus, the generated code is expected to perform well across multiple datasets.
As shown in Figure 22, parent 1 and parent 2 excel on dataset 4 and dataset 3, respectively. By combining their complementary strengths through the mixing strategy performed by the LLM in Table 4, the offspring operator achieves strong performance on both datasets 3 and 4, thereby yielding better average performance. Unlike random subtree crossover in tree-based genetic programming (Banzhaf et al. 1998) or one-point crossover in linear genetic programming (Huang et al. 2024), LLM-based crossover can synthesize code that semantically resembles an intermediate between the two parent programs by leveraging its code understanding capability. This highlights the importance of semantic diversity: if both parents are highly similar, even if performant, the resulting offspring would be nearly identical to the parents, leading to redundant evaluations and wasted computation. Therefore, semantic-aware evolution is vital for automated algorithm design.
Figure 22: Validation $R^{2}$ scores of parent and offspring selection operators on different datasets.
Component Parent 1 Parent 2 Generated Code
Complexity penalty $(s+h)\times\dfrac{0.3+0.7\,\text{stage}}{100}$ $\dfrac{s+h}{30+70\,\text{stage}}$ $(s+h)\times\dfrac{0.4+0.6\,\text{stage}}{40+60\,\text{stage}}$
Subset size $\max\!\bigl{(}7,\lfloor\tfrac{n_{c}}{\tfrac{k}{2}+3}\rfloor\bigr{)}$ $\max\!\bigl{(}10,\lfloor\tfrac{n_{c}}{\tfrac{k}{2}+1}\rfloor\bigr{)}$ $\max\!\bigl{(}8,\lfloor\tfrac{n_{c}}{\tfrac{k}{2}+2}\rfloor\bigr{)}$
Parent A scoring $(1-\text{stage})\times(\mathrm{mse}_{\text{full}}-\mathrm{mse}_{\text{sub}})\;+\;\text{stage}\times\dfrac{1}{\mathrm{mse}_{\text{full}}}\;-\;\mathrm{comp\_pen}$ $(1-\text{stage})\times\dfrac{1}{\mathrm{mse}_{\text{sub}}}+\;\text{stage}\times\dfrac{1}{\mathrm{mse}_{\text{full}}}-\;\mathrm{comp\_pen}$ Same as Parent 1
Parent B scoring $r+\dfrac{\mathrm{comp\_pen}}{1+5\,\text{stage}} corr+co$$r+\dfrac{\mathrm{comp\_pen}}{10}$ $corr+\dfrac{\mathrm{comp\_pen}}{\begin{cases}5,&\text{if stage}<0.5\\ 10,&\text{otherwise}\end{cases}}$
Table 4: Fusion of key selection components. Here $s=\text{sizes}$, $h=\text{heights}$, $n_{c}$ is the number of cases, $\mathrm{comp\_pen}$ is the complexity penalty, and corr denotes the normalized residual correlation.
1def omni_selection(population, k=100, status={}):
2 stage = status.get("evolutionary_stage", 0)
3 n = len(population)
4 half_k = k // 2
5 rng = np.random.default_rng(12345)
6
7 errs = np.array([ind.case_values for ind in population]) # (n,n_cases)
9 sizes = np.array([len(ind) for ind in population])
10 heights = np.array([ind.height for ind in population])
11
12 n_cases = errs.shape[1]
13 comp_pen = (sizes + heights) * (0.4 + 0.6 * stage) / (40 + 60 * stage) # adaptive comp penalty
14
15 subset_size = max(8, n_cases // (half_k + 2))
16 max_tries = 15 * half_k
17 tried_subsets = set()
18 full_mse = errs.mean(axis=1)
19
20 parent_a = []
21 tries = 0
22 while len(parent_a) < half_k and tries < max_tries:
23 tries += 1
24 subset = tuple(sorted(rng.choice(n_cases, subset_size, replace=False)))
25 if subset in tried_subsets:
26 continue
27 tried_subsets.add(subset)
28
29 mse_sub = errs[:, subset].mean(axis=1)
30 specialization = full_mse - mse_sub # + means better on subset specialization
31 scores = (
32 (1 - stage) * specialization +
33 stage * (1 / (full_mse + 1e-10))
34 ) - comp_pen
35 best_idx = np.argmax(scores)
36 parent_a.append(population[best_idx])
37
38 if not parent_a:
39 parent_a = rng.choice(population, half_k, replace=True).tolist()
40
41 norm_resid = residuals / (np.linalg.norm(residuals, axis=1, keepdims=True) + 1e-10)
42 parent_b = []
43 comp_factor = 5 if stage < 0.5 else 10 # stronger complexity pressure late
44 for a in parent_a:
45 a_idx = population.index(a)
46 corr = norm_resid @ norm_resid[a_idx]
47 corr[a_idx] = 1 # exclude self
48 scores = corr + comp_pen / comp_factor
49 comp_idx = np.argmin(scores)
50 parent_b.append(population[comp_idx])
51
52 selected = [ind for pair in zip(parent_a, parent_b) for ind in pair]
53 return selected[:k]
Code 4: Offspring Code
1def omni_selection(population, k=100, status={}):
2 stage = status.get("evolutionary_stage", 0)
3 n = len(population)
4 half_k = k // 2
5
6 errs = np.array([ind.case_values for ind in population]) # (n, n_cases)
8 sizes = np.array([len(ind) for ind in population])
9 heights = np.array([ind.height for ind in population])
10 n_cases = errs.shape[1]
11 rng = np.random.default_rng(777)
12
13 # Complexity penalty grows with stage, stronger late to favor interpretability
14 comp_pen = (sizes + heights) * (0.3 + 0.7 * stage) / 100
15
16 # Parameters for subset selection
17 subset_size = max(7, n_cases // (half_k + 3))
18 max_tries = 10 * half_k
19 tried_subsets = set()
20 full_mse = errs.mean(axis=1)
21
22 parent_a = []
23 while len(parent_a) < half_k and max_tries > 0:
24 max_tries -= 1
25 subset = tuple(sorted(rng.choice(n_cases, subset_size, replace=False)))
26 if subset in tried_subsets:
27 continue
28 tried_subsets.add(subset)
29
30 mse_sub = errs[:, subset].mean(axis=1)
31 specialization = full_mse - mse_sub # positive means better on subset
32 # Blend specialization and overall fitness with stage-dependent weights
33 scores = (1 - stage) * specialization + stage * (1 / (full_mse + 1e-10))
34 scores -= comp_pen
35 best_idx = np.argmax(scores)
36 parent_a.append(population[best_idx])
37
38 if not parent_a:
39 parent_a = rng.choice(population, half_k, replace=True).tolist()
40
41 norm_resid = residuals / (np.linalg.norm(residuals, axis=1, keepdims=True) + 1e-10)
42 parent_b = []
43 for a in parent_a:
44 a_idx = population.index(a)
45 corr = norm_resid @ norm_resid[a_idx]
46 corr[a_idx] = 1
47 # Score complementary: minimize correlation and complexity; complexity effect softer early on
48 scores = corr + comp_pen / (1 + stage * 5)
49 b_idx = np.argmin(scores)
50 parent_b.append(population[b_idx])
51
52 selected = [ind for pair in zip(parent_a, parent_b) for ind in pair]
53 return selected[:k]
Code 5: Code of Parent A
1def omni_selection(population, k=100, status={}):
2 n = len(population)
3 half_k = k // 2
4 stage = status.get("evolutionary_stage", 0)
5 max_tries = 10 * k
6
7 preds = np.array([ind.predicted_values for ind in population])
8 cases = np.array([ind.case_values for ind in population])
10 sizes = np.array([len(ind) for ind in population])
11 heights = np.array([ind.height for ind in population])
12
13 n_cases = preds.shape[1]
14 subset_size = max(10, n_cases // (half_k + 1))
15 idx_pool = np.arange(n_cases)
16 np.random.seed(42)
17
18 # Complexity penalty scaled by stage
19 comp_pen = (sizes + heights) / (30 + 70 * stage)
20
21 # Score function balances subset/full MSE and complexity
22 def score(i, subset_idx):
23 mse_subset = np.mean((cases[i, subset_idx] - preds[i, subset_idx]) ** 2)
24 mse_full = np.mean((cases[i] - preds[i]) ** 2)
25 return ((1 - stage) * (1 / (mse_subset + 1e-10)) + stage * (1 / (mse_full + 1e-10))) - comp_pen[i]
26
27 parent_a = []
28 tried_subsets = set()
29 tries = 0
30 while len(parent_a) < half_k and tries < max_tries:
31 tries += 1
32 subset_idx = tuple(sorted(np.random.choice(idx_pool, subset_size, replace=False)))
33 if subset_idx in tried_subsets:
34 continue
35 tried_subsets.add(subset_idx)
36 scores = np.array([score(i, subset_idx) for i in range(n)])
37 best_idx = np.argmax(scores)
38 parent_a.append(population[best_idx])
39
40 if not parent_a:
41 parent_a = np.random.choice(population, half_k, replace=True).tolist()
42
43 parent_b = []
44 norm_residuals = residuals / (np.linalg.norm(residuals, axis=1, keepdims=True) + 1e-10)
45 for a in parent_a:
46 a_idx = population.index(a)
47 a_res = norm_residuals[a_idx]
48 corr = norm_residuals @ a_res
49 corr[a_idx] = 1.0 # exclude self
50 complement_scores = corr + comp_pen / 10 # prefer low corr & low complexity
51 comp_idx = np.argmin(complement_scores)
52 parent_b.append(population[comp_idx])
53
54 selected_individuals = [ind for pair in zip(parent_a, parent_b) for ind in pair]
55 return selected_individuals[:k]
Code 6: Code of Parent B
Appendix E Results on More Discovered Operators
The proposed LLM-Meta-SR method can discover various selection operators. For example, LABEL:lst:offspring presents a selection operator discovered by the LLM-Meta-SR algorithm, which differs from the operator shown in LABEL:lst:omni-selection. To avoid confusion, we refer to the discovered operator in LABEL:lst:offspring as “Holo”. The results of the test $R^{2}$ scores are shown in Figure 24 and Figure 24, which demonstrate that the Holo selection operator outperforms expert-designed selection operators. Additionally, the complexity size, shown in Figure 25, indicates that the discovered operator is competitive in terms of model complexity. A direct comparison with the tree sizes evolved using the Omni selection, reported in Figure 10, shows that the discovered operator is even slightly smaller. Therefore, the “Holo” selection operator is preferred when interpretability is a key concern.
Figure 23: Test $R^{2}$ scores of different selection operators on symbolic regression benchmarks, including the Holo selection operator.
Figure 25: Tree sizes of different selection operators on symbolic regression benchmarks, including the Holo selection operator.
Appendix F Test Cases for Synthetic Evaluation
@@
Appendix G More Analysis on Symbolic Regression Benchmark
The results of the top-5 performing algorithms are shown in Figure 26 for better clarity, and the Pareto front of the ranks of test $R^{2}$ scores and model sizes across different algorithms is shown in Figure 27. These results demonstrate that RAG-SR-Omni is a Pareto-optimal algorithm on SRBench. Specifically, RAG-SR-Omni dominates retrieval-augmentation-generation-based symbolic regression (RAG-SR) (Zhang et al. 2025) as well as transformer-based planning for symbolic regression (Shojaee et al. 2024). Overall, these results show the effectiveness of the LLM-evolved selection operator.
Figure 27: Pareto front of the ranks of test $R^{2}$ scores and model sizes for different algorithms.
Appendix H Baseline Selection Operators
The hyperparameters for the baseline selection operators follow the default settings specified in their respective original papers. For RDS-Tour, the sampling ratio is set to 10% and the tournament size to 7 (Geiger et al. 2025). For PLex selection, the temperature parameter is set to 1.0 (Ding, Pantridge, and Spector 2023). For DALex, a particularity pressure of 3 is used (Ni, Ding, and Spector 2024). For CPS, the first parent is selected via tournament, with the tournament size set to 7 (Xu et al. 2022).
For Boltzmann sampling with temperature scheduling (Shojaee et al. 2025; Romera-Paredes et al. 2024), a temperature decay rule is applied to balance exploration and exploitation over time, similar to the approaches used in LLM-SR (Shojaee et al. 2025) and FunSearch (Romera-Paredes et al. 2024). The sampling probability for individual $i$ is defined as:
@@ Appendix H Baseline Selection Operators
where $s_{i}$ denotes the score of individual $i$, $\tau_{0}=0.1$ is the initial temperature (Shojaee et al. 2025), $t$ is the current generation, and $T_{\max}$ is the total number of generations. This scheduling rule linearly decays the temperature over generations and gradually shifts the selection pressure from exploration (high temperature) to exploitation (low temperature) as evolution progresses in evolutionary symbolic regression.
Appendix I Further Investigation on Crossover and Mutation Ratio
@@
In recent years, lexicase selection has shown strong empirical performance and is widely adopted in modern symbolic regression systems (Cava et al. 2019; Zhang et al. 2025). Inspired by lexicase selection, several variants have been developed, including probabilistic lexicase selection (Ding, Pantridge, and Spector 2023), lexicase-like selection via diverse aggregation (DALex) (Ni, Ding, and Spector 2024), $\epsilon$-lexicase selection with dynamic split (D-Split) (Imai Aldeia, De França, and La Cava 2024), down-sampled lexicase selection (Boldi et al. 2024), and random down-sampled tournament selection (Geiger, Sobania, and Rothlauf 2025). These methods share the core idea of emphasizing performance on subsets of training instances to encourage specialization, and they demonstrate advantages in various scenarios (Geiger, Sobania, and Rothlauf 2025). However, these operators are manually designed and do not fully satisfy all the desired properties outlined in Section 3.5, highlighting the need for further exploration of selection operators.
@@
Before the LLM era, genetic programming had been widely used for automated evolutionary algorithm design, including the design of fitness functions (Fong and Motani 2024a), parameter adaptation techniques (Stanovov, Akhmedova, and Semenkin 2022), and update strategies (Lones 2021). More recently, reinforcement learning has also been applied to designing update rules in differential evolution (Chen et al. 2024a). However, these approaches are typically limited to generating only small code segments, as exploring large program spaces remains challenging. The emergence of LLMs has made automated algorithm design more feasible. For example, LLMs have been used to design a differential evolution algorithm competitive with state-of-the-art continuous optimization methods (Stein and Bäck 2024). Similarly, LLMs have been employed to improve the self-organizing migrating algorithm (Pluhacek et al. 2024). This growing interest in LLMs motivates the development of efficient LLM-based approaches for automated algorithm design, where semantic awareness and code bloat are two key issues that remain underexplored and warrant further investigation.
@@
LLMs have been employed as optimization tools across various domains (Meyerson et al. 2024; Song et al. 2024), including neural architecture search (Chen, Dohan, and So 2023), graph neural network design (Mo et al. 2025), large neighborhood search algorithm design (Ye et al. 2025), reward function design (Ma et al. 2024), the traveling salesman problem (Liu et al. 2024b), prompt optimization (Yang et al. 2024), automated machine learning (Chan et al. 2025), solver code generation (Jiang et al. 2025), symbolic regression (Shojaee et al. 2025), and feature engineering (Gong et al. 2025). Evolutionary search has gained considerable attention in the LLM era due to its effectiveness over random sampling (Zhang et al. 2024). Various efforts have aimed to enhance LLM-based evolutionary search, including group relative policy optimization (Huang et al. 2025b), direct preference optimization (Šurina et al. 2025), core abstraction prompting (Wu et al. 2025), quality-uncertainty optimization (Chen et al. 2024b), and bi-dimensional co-evolution (Zhao et al. 2025).